OpenStudy (yellowlegoguy99):

Which expression is equivalent to 16^6/16^2 16^3 16^4 16^-4 16^-3

OpenStudy (yellowlegoguy99):

@DSS @dg98

OpenStudy (anonymous):

65536

OpenStudy (yellowlegoguy99):

...?

OpenStudy (anonymous):

Therefore your answer is 16^4

OpenStudy (yellowlegoguy99):

How do you do these so fast? o.O

OpenStudy (anonymous):

I agree with you Night:)) Your answer is 16^4..Hope it helps:))

OpenStudy (anonymous):

when dividing exponents with like terms, you subtract exponents. so 16^6/16^2 =16^6-2 =16^4